OpenStudy (anonymous):

prove that sin20 sin40 sin60 sin80=3/16

OpenStudy (shubhamsrg):

What is your attempt so far?

OpenStudy (anonymous):

calculator.lol dats all you need

OpenStudy (anonymous):

Hint: \[sinx.\sin(60-x).(\sin60+x) = \frac{ \sin 3x }{ 4 }\]

OpenStudy (anonymous):

sin60 is \[\sqrt{3}/2\] right? put it there in that, and then apply the above formula you'll get it... \[\sqrt{3}/2\] * sin60/4 = 3/2*4*2 = 3/16